Why transformer coupling is not
preferred at low frequencies?
1
Expert's answer
2019-02-21T11:18:50-0500

Transformers are constructed to transform voltage or (sometimes) current. Let's look at the transformer EMF equation:


"E=\frac{2\pi}{\sqrt{2}}\Phi f w=4.44\Phi f w,"


where:

"\Phi" - the magnetic flux on the core, it depends on cross-section of the core and the magnetic field "B" inside, which is limited by approximately 1.5 teslas; thus high values of "\Phi" are limited by properties of material, amount of material used, and cooling conditions.

"w" - number of turns in the transformer's winding, limited by amount of material used and cooling conditions.

"f" - frequency. Roughly saying unlimited value.

That is why to achieve the required EMF it's much more effective to increase frequency than the other parameters.